The sum of 3 consecutive integers is -60 what’s the integers
1
Expert's answer
2020-11-05T13:25:01-0500

x + (x+1) + (x+2) = -60

3x + 3 = -60

3x = -60-3 = -63

x = (-63/3) = -21

x+1 = -21+1 = -20

x+2 = -21+2 = -19


Three numbers are -19, -20, -21